OpenStudy (anonymous):

if x is an integer, which of the following could be x^3 a) 2.7 x 10^11 b)2.7 x 10^12 c)2.7 x 10^13 d)2.7 x 10^14 e)2.7 x 10^15

OpenStudy (anonymous):

can you express any of these numbers as a cube of an integer?

OpenStudy (anonymous):

hmm 2.7x10^13=27x10^12=(3x10^4)^3=30000^3 so its c and x=30000